Two forces of 80N and 100N acting at an angle of 60°with each other pull an object. Determine the net force using the parallelogram method.  


1
Expert's answer
2021-01-04T14:33:03-0500

"F=\sqrt{80^2+100^2+2\cdot80\cdot100\cdot\cos60°}=156.2(N)"